Marigold Corp. purchased a truck for $66,000. The company expected the truck to last four years or 102,000 miles, with an estimated residual value of $15,000 at the end of that time. During the second year the truck was driven 47,000 miles. Compute the depreciation for the second year under each of the methods below and place your answers in the blanks provided.

Units-of-activity $enter the depreciation for the second year under the Units-of-activity method in dollars

Double-declining-balance $enter the depreciation for the second year under the Double-declining-balance method in dollars
Answer #1

Double declining rate = (66,000-15,000)/102,000 = 0.5 per mile

Depreciation for second year = 0.5 * 47,000

= 23,500

Double declining rate = 200/Useful life = 200/4 = 50%

Depreciation for first year = 66,000*50% = 33,000

Depreciation for second year = (66,000-33,000)*50%

= 16,500

Comment if you face any issues